# BALF: Budgeted Activation-Aware Low-Rank Factorization for Fine-Tuning-Free Model Compression

### v1

@@ Appendix I A Post-Hoc Interpretation via Information Geometry
$\bigl\|f({\bm{\mathsfit{X}}};{\bm{\mathsfit{W}}})-f\bigl({\bm{\mathsfit{X}}};\mathcal{T}^{\mathrm{AA}}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)\bigr)\bigr\|_{F}$ $=\bigl\|\overline{O}\!\bigl(\overline{I}({\bm{\mathsfit{X}}})\,\overline{P}({\bm{\mathsfit{W}}})\bigr)-\overline{O}\!\bigl(\overline{I}({\bm{\mathsfit{X}}})\,\mathcal{T}^{\mathrm{AA}}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\bigr)\bigr\|_{F}$ (3)
@@
$=\frac{1}{B}\left\|\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}\left({\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})-\mathcal{T}^{\mathrm{SVD}}_{P}\bigl({\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})\bigr)\right)\right\|_{F}^{2}$
@@
$\underset{(a)}{=}\frac{N}{B}\operatorname{tr}\left({\bm{V}}{\bm{\Sigma}}^{2}{\bm{V}}^{T}\right)$
@@
For step (a), note the following. First, recall from the definition of $\bm{M}$} (see Section 3) that ${\bm{M}}^{T}\overline{I}({\bm{\mathsfit{X}}})^{T}\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}=N\overline{{\bm{I}}}_{R}$ (where $R$ is the rank of $\overline{I}({\bm{\mathsfit{X}}})$); and that $\bm{M}$} has its last IRI-R columns set to 0, meaning $\bm{M}$$^{+}$ has its last IRI-R rows set to 0.
This results in ${\bm{U}}^{T}{\bm{M}}^{T}\overline{I}({\bm{\mathsfit{X}}})^{T}\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}{\bm{U}}=N{\bm{U}}^{T}\overline{{\bm{I}}}_{R}{\bm{U}}=N\overline{{\bm{I}}}_{R}$ (this last equality follows from the fact that ${\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})\in\mathbb{R}^{I\times O}$ has its last IRI-R rows set to 0 since $\bm{M}$$^{+}$ has).
@@
Let ${\bm{Y}}\in\mathbb{R}^{M\times N}$ satisfy ${\bm{Y}}^{T}{\bm{Y}}=\alpha{\bm{I}}_{N}$ with $\alpha>0$. Then for any ${\bm{X}}\in\mathbb{R}^{N\times K}$ and any $P\geq 1$,
@@
Hence the (nonzero) eigenvalues of $({\bm{Y}}{\bm{X}})^{T}({\bm{Y}}{\bm{X}})$ are exactly $\alpha$ times those of ${\bm{X}}^{T}{\bm{X}}$, so $\bm{\sigma}_{i}({\bm{Y}}{\bm{X}})=\sqrt{\alpha}\bm{\sigma}_{i}({\bm{X}})$ for all $i$. Summing the top $P$ singular values gives the claim. ∎
@@
We need to prove that the sum of singular values up to $P$ of both quantities is equivalent (see Definition 2 for context). It suffices to prove the result for a single group. Fix a group (omitted from the notation for simplicity) and let ${\bm{W}}=\overline{P}({\bm{\mathsfit{W}}})\in\mathbb{R}^{I\times O},{\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})\in\mathbb{R}^{N\times I}$, ${\bm{M}}\in\mathbb{R}^{I\times I}$, ${\bm{M}}^{+}\in\mathbb{R}^{I\times I}$ denote the items corresponding to said group, and denote the rank of $\bm{X}$} by $R$.
By construction, $\bm{M}$} has its last IRI-R columns set to 0 (see Section 3), which forces $\bm{M}$$^{+}$ to have its last IRI-R rows set to 0. Hence, ${\bm{M}}{\bm{M}}^{+}={\bm{M}}{\bm{I}}_{:,:R}{\bm{I}}_{:R,:}{\bm{M}}^{+}$.
Moreover, from the definition of $\bm{M}$} we have ${\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}=N\overline{{\bm{I}}}_{R}$, and therefore ${\bm{I}}_{:R,:}{\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}{\bm{I}}_{:,:R}=N{\bm{I}}_{R}$.
It is also true that ${\bm{X}}{\bm{M}}{\bm{M}}^{+}={\bm{X}}$ (see Section 3).
@@
where the last equality follows from Lemma 1. The result now follows from Theorem 2.
Appendix K LLM Usage
@@
LLMs were used to bootstrap, modify, and evaluate code (including checking correctness). They were particularly useful for iterating on result-visualization scripts (e.g., code that takes experimental outputs and produces plots or table components).

### v2

@@ I.1 Proof of Theorem 2
$\bigl\|f({\bm{\mathsfit{X}}};{\bm{\mathsfit{W}}})-f\bigl({\bm{\mathsfit{X}}};\mathcal{T}^{\mathrm{AA}}_{P}\bigl({\bm{\mathsfit{W}}}\bigr)\bigr)\bigr\|_{F}$ $=\bigl\|\overline{O}\!\bigl(\overline{I}({\bm{\mathsfit{X}}})\,\overline{P}({\bm{\mathsfit{W}}})\bigr)-\overline{O}\!\bigl(\overline{I}({\bm{\mathsfit{X}}})\,\mathcal{T}^{\mathrm{AA}}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\bigr)\bigr\|_{F}$
@@
$\underset{(a)}{=}\frac{1}{B}\left\|\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}\left({\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})-\mathcal{T}^{\mathrm{SVD}}_{P}\bigl({\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})\bigr)\right)\right\|_{F}^{2}$
@@
$\underset{(b)}{=}\frac{N}{B}\operatorname{tr}\left({\bm{V}}{\bm{\Sigma}}^{2}{\bm{V}}^{T}\right)$
@@
For step (a), note that $\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}{\bm{M}}^{+}=\overline{I}({\bm{\mathsfit{X}}})$ (see Section 3).
For step (b), note the following. First, recall from the definition of $\bm{M}$} (see Section 3) that ${\bm{M}}^{T}\overline{I}({\bm{\mathsfit{X}}})^{T}\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}=N\overline{{\bm{I}}}_{R}$ (where $R$ is the rank of $\overline{I}({\bm{\mathsfit{X}}})$); and that $\bm{M}$} has its last IRI-R columns set to 0, meaning $\bm{M}$$^{+}$ has its last IRI-R rows set to 0.
This results in ${\bm{U}}^{T}{\bm{M}}^{T}\overline{I}({\bm{\mathsfit{X}}})^{T}\overline{I}({\bm{\mathsfit{X}}}){\bm{M}}{\bm{U}}=N{\bm{U}}^{T}\overline{{\bm{I}}}_{R}{\bm{U}}=N\overline{{\bm{I}}}_{R}$ (this last equality follows from the fact that ${\bm{M}}^{+}\overline{P}({\bm{\mathsfit{W}}})\in\mathbb{R}^{I\times O}$ has its last IRI-R rows set to 0 since $\bm{M}$$^{+}$ does).
@@ I.2 Proof of Theorem 1
Let ${\bm{Y}}\in\mathbb{R}^{M\times N}$ satisfy ${\bm{Y}}^{T}{\bm{Y}}=\alpha{\bm{I}}_{N}$ with $\alpha>0$. Then, for any ${\bm{X}}\in\mathbb{R}^{N\times K}$ and any $P\geq 1$,
@@
Hence, the (nonzero) eigenvalues of $({\bm{Y}}{\bm{X}})^{T}({\bm{Y}}{\bm{X}})$ are exactly $\alpha$ times those of ${\bm{X}}^{T}{\bm{X}}$, so $\bm{\sigma}_{i}({\bm{Y}}{\bm{X}})=\sqrt{\alpha}\bm{\sigma}_{i}({\bm{X}})$ for all $i$. Summing the top $P$ singular values gives the claim. ∎
@@
It suffices to prove the result for a single group. Fix a group (its index is omitted from the notation for brevity), and let ${\bm{W}}=\overline{P}({\bm{\mathsfit{W}}})\in\mathbb{R}^{I\times O},{\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})\in\mathbb{R}^{N\times I}$, ${\bm{M}}\in\mathbb{R}^{I\times I}$, and ${\bm{M}}^{+}\in\mathbb{R}^{I\times I}$ denote the items corresponding to said group, and denote the rank of $\bm{X}$} by $R$.
We recall the following facts; see Section 3. First, $\bm{M}$} has its last IRI-R columns set to 0, which forces $\bm{M}$$^{+}$ to have its last IRI-R rows set to 0. Hence, ${\bm{M}}{\bm{M}}^{+}={\bm{M}}{\bm{I}}_{:,:R}{\bm{I}}_{:R,:}{\bm{M}}^{+}$. Moreover, from the definition of $\bm{M}$}, we have ${\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}=N\overline{{\bm{I}}}_{R}$, and therefore, it holds that ${\bm{I}}_{:R,:}{\bm{M}}^{T}{\bm{X}}^{T}{\bm{X}}{\bm{M}}{\bm{I}}_{:,:R}=N{\bm{I}}_{R}$. It is also true that ${\bm{X}}{\bm{M}}{\bm{M}}^{+}={\bm{X}}$.
It follows that
where the last equality follows from Lemma 1.
@@
$\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\mathcal{T}^{\mathrm{SVD}}_{P}\bigl(\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})\bigr)\|^{2}_{F}$ $=\sum_{i=P+1}^{U}\bm{\sigma}_{i}^{2}({\bm{X}}{\bm{W}})$
$=N\sum_{i=P+1}^{U}\bm{\sigma}_{i}^{2}({\bm{M}}^{+}{\bm{W}})$
$=\|\overline{I}({\bm{\mathsfit{X}}})\overline{P}({\bm{\mathsfit{W}}})-\overline{I}({\bm{\mathsfit{X}}})\mathcal{T}^{\mathrm{AA}}_{P}\bigl(\overline{P}({\bm{\mathsfit{W}}})\bigr)\|^{2}_{F},$
where the last equality follows from Theorem 2.
Taking square roots completes the proof. ∎
Appendix J LLM Usage
@@
LLMs were used to bootstrap, modify, and evaluate code (including checking for correctness). They were particularly useful for iterating on result-visualization scripts (e.g., code that takes experimental outputs and produces plots or table components).